Solve using the elimination method.
-5x - 7y = 25.5
-2x + 10y = -9

Answer by sofiyac(983)

-5x - 7y = 25.5 multiply first one by 2, second one by -5 and add them
-2x + 10y = -9
***
-10x-14y=51
10x-50y=45
------------
-64y=96 divide each side by -64
y=-1.5 plug it into one of the original equations and solve for x
-2x + 10y = -9
-2x + 10*(-1.5) = -9
-2x-15=-9
-2x=6 divide each side by -2
x=-3
solution is (-3,-1.5)